# Surveying optically addressable spin qubits for quantum information and sensing technology

### v2

@@ Abstract
Keywords spin qubits quantum sensing spintronics materials engineering
1 Introduction
@@
To be useful, spintronic materials must enable both single-spin addressability and long storage times at high levels of fidelity. These properties are, in turn, limited by the materials’ spin-lattice relaxation () and spin decoherence times (, also known as phase memory time). and represent the time it takes for the "longitudinal" and "transverse" magnitude of the spin state vector, respectively, to decay by a factor of e. In practical terms, is the lifetime of a non-Boltzmann population (Figure 1b), and, put simply by DiVincenzo [8], characterises the interactions of a qubit with its environment. These interactions are ultimately responsible for causing a qubit in a generic (pure) state to relax to a mixed state, leading to the loss of quantum information via phase decoherence (Figure 1c). These parameters can be measured using electron spin resonance (ESR) techniques, which measure light/microwave absorptions and emissions from a material as a function of frequency or magnetic field strength (Figure 2). is typically measured by Hahn-Echo pulses but can also be determined from the homogeneous emission linewidth of its resonance frequency, (Figure 1d). In practice, spin systems exhibit an inhomogeneously broadened linewidth, , due to differences in the magnetic field, B, at the locations of otherwise equivalent spins (see Figure 1e). These differences are in turn due to differences in the locations and orientations of the electronic and/or nuclear spins that surround each measured spin. Following Chavhan et al., the relationship between and can be represented as 1/ = 1/ + B, where is the gyromagnetic ratio [9]. In general, sets an upper limit on and, since energy relaxation proceeds more slowly than decoherence, is (often substantially) longer than [10]. Likewise, sets an upper limit on (the pulse sequence used to measure deliberately attempts to reduce the effects of inhomogeneity). Thus, .
@@
Quantum systems capable of operating at room temperature open up many additional applications that the overhead of cryogenic operation precludes. But, achieving high fidelity (in the initialisation, gate operations and read-out) of room-temperature qubits remains extremely challenging. In this review, we have attempted to survey the available "fully optically addressable materials" (FOAMs) that are most directly relevant to quantum information and sensing technology [11]. We do not provide an exhaustive account of all reported spin systems. Rather, we present the different approaches, materials-wise, that have shown promise and report the properties (where available) of the best-performing representatives of each approach.
2 Overview of Candidate Spin Qubit Materials
@@
FOAMs are an attractive option for quantum applications since they typically involve spin levels that are energetically separated by far more than and can be initialised using visible/near-infrared light to generate quasi-"pure" quantum states. Furthermore, optical signals, being themselves far larger than , are less affected by thermal noise and benefit from excellent single-photon detectors that are available even at room temperature. Since the energies of the photons involved in both processes greatly exceed , room-temperature operation at high levels of fidelity is possible. Such materials are usually probed using ODMR spectroscopy, where the quantum sensing sensitivity for a.c. magnetic fields (for example), , is proportional to , where is the duration of the readout process, is the measurement contrast, is the number of spins and is the number of scan averages. Relatively few FOAMs have been demonstrated so far and are often limited by either low photoluminescence yields or relatively short-lived spin polarizations () and coherences (), which have been summarised in Figure 3 to support our discussion. As a prerequisite for inclusion in figure 3, we have opted to follow the criteria put forward by Weber et al., for the realization of an optically addressable solid-state spin qubit (that is useable)[12]:
@@
2.1 3D spin-defect FOAMs
2.1.1 Diamond spin centres
A well-established example of a FOAM which satisfies the above criteria is the negatively charged nitrogen-vacancy centres in diamond (herein simply, NV-diamond), which utilise the triplet ground state and hyperfine splitting arising from the NV-centre. Initialisation (i.e., electron spin polarisation) is achieved by optical excitation with green light. The newly generated excited states either relax by emission of 637 nm light or undergo intersystem crossing (ISC) into a metastable singlet state (Figure 4a). Repopulation of the ground state then follows an intermediate relaxation between two singlet states (with emission at 1042 nm) and finally spin-selective ISC into the T0 sublevel, resulting in a strong spin polarisation. At zero-applied magnetic field (ZF) the T±1 states are degenerate due to the defects C3V symmetry. The application of a magnetic field lifts this degeneracy, which enables spin manipulation using microwave pulses, thereby modifying the fluorescence at characteristic Zeeman splitting frequencies (), which can be detected following subsequent optical excitations and fluorescence back to T0.
NV-diamond has been a cornerstone of ODMR-based quantum sensing due to its robust spin properties even at room temperature. Due to the relatively low spin densities (resulting in small dipolar coupling between spins) and the mismatch between diamond lattice vibrations (phonons) and the Larmor frequency of electron spins in NV-’s S=1 ground state, NV-diamond can exhibit s of several milliseconds at room temperature [60, 61]. The nature and mechanism of its spin-lattice relaxation as a function of temperature, which includes an ‘Orbach-like" process (dependent on the phonon density at the spin transition frequency) and spin-phonon Raman scattering (of either first [62] or else second [61] order) scaling like , have been repeatedly investigated.
@@ 2.1.1 Diamond spin centres
2.1.2 Silicon Carbide Spin Centres
@@
Beyond diamond, silicon carbide (SiC) also shows promising optical and coherence properties for quantum applications [85, 86, 87, 88, 89, 90]. SiC is a complex material with over 200 polymorphs. It is also used commercially in electronics and hence benefits from decades of manufacturing experience. Mercifully, the study of quantum systems has largely been restricted to 3C-, 4H-, and 6H-SiC, where C and H signify the cubic and hexagonal structures, respectively, and the preceding number designates its polytype [86] (see Figure 5a). Pure SiC has a wide bandgap (2 - 3 eV), weak spin-orbit coupling, and a naturally low abundance of nuclear spins. Importantly, it is capable of harbouring several different optically addressable colour centres with (often) near-infrared emission and record ODMR contrasts[31]. The most commonly studied defects consist of neutral or negatively-charged monovacancies (VC or VSi) and divacancies (VV) but can also include carbon anti-site vacancies (CSiVC) (see Figure 5b and c), charged NVs, Cr4+- [41, 91], V3+, V4+ [92], Mo5+ [44], and Ti-centres [93, 86, 90]. Further layers of complexity are added by consideration of additional oxidation states and crystallographic sites within each polytype. Understandably, studies often focus on the "best performing" defect within a given sample, and seldom are single samples of SiC homogeneous. This represents the material’s most significant practical limitation as a spin qubit candidate compared to other platforms.
@@
Nevertheless, several defects exhibit outstanding spin-optical properties, and the fabrication challenges are beginning to be addressed. Negatively charged VSi in isotopically purified (28Si) 4H-SiC exhibits the largest ODMR contrast at 97 at 4 K [31]. With an S=3/2 ground state, its associated ZFS is a few MHz with degenerate pairs of |1/2 and |3/2 states (Figure 5d). Under a magnetic field (82 mT) precisely aligned to the crystallographic c-axis, this degeneracy is lost and in the excited, so-called "V1" state, |3/2 shifts higher energy than the |1/2 states. To achieve 97 contrast, the authors first equilibrate the spin populations using a 40 s off-resonance pump (at 730 nm), followed by an on-resonance pump (at 861 nm) lasting up to 80 s. On-resonance optical pumping causes |3/2 states to selectively decay into a non-radiative metastable state, followed by spin-selective repopulation of |1/2 ground states. This leads to a spin polarisation of up to 90 which is paired with record s and s of 0.80.12 ms and 302 s, respectively. More recently, it was shown that VSi defects can be implanted into nanophotonic waveguides fabricated from 4H-SiC while also controlling the alignment of individual defects and maintaining excellent spin-optical properties [32]. Here, of bulk VSi-centres was measured at 344 s at 10 K, whilst those in the 1 m diameter waveguides measured at 9.40.7 s. The coherence properties can be further improved by a factor of 10 through a regime of isotopic purification, and another factor of 5 by reducing strain inhomogeneity through a regime of annealing. Using this combined approach, Lekavicius et al., enhanced from 400 ns to 20 s at room-temperature [33].
@@
2.1.3 Inorganic Semiconductor Spin Centres
Beyond vacancy systems, several metal-ion centres have demonstrated interesting spin-optical properties. By far the best, however, is Cr4+ in 4H-SiC. As a S=1 species, this material exhibits >1 second, and and = 81 s and 317 ns, respectively, at 15 K [91]. Importantly, it also demonstrates a 79 contrast, marking it as a system with some of the highest optical readout fidelity. The spin-optical properties are markedly impaired in GaN host, which demonstrates 27x broader emission linewidths due to interactions of Cr4+-centres with the surrounding spin bath [41].
2.2 2D- and van der Waals FOAMs
@@ 2.1.3 Inorganic Semiconductor Spin Centres
The prototypical vdW material, graphene, was an early candidate for harbouring spin qubits; however, the metallic band structure of graphene is not conducive to forming luminescent spin centres. Instead, its isoelectronic analogue, hBN, has been the centre of remarkable advancement as a 2D-optically addressable platform at room temperature, building off its success as a single-photon source [103]. Magnetically-active spin centres in hBN were identified by EPR in the 1960s, with hyperfine structures giving clues to their identity as S=1/2 boron centres and carbon impurities [104, 105, 106]. However, the first optically addressable spin centres were only identified recently as triplets stemming from a negatively charged vacancy, likely originating from a missing boron atom, so-called -centres [107, 108, 109, 110] (Figure 6a). Similarly to NV-diamond centres, optical initialisation is achieved by photoexcitation with 532 nm light, leading to ISC into a metastable singlet, which spin-selectivity repopulates the fluorescent ms=0 state (TZ at zero-applied field, Figure 6b). The readout of the spin state occurs through 850 nm fluorescence with pulsed ODMR contrast of 1 at room temperature and almost 30 at 10 K, revealing a ZFS of |D|3.5GHz (Figure6c). Spin-relaxation and Hahn-echo measurements reveal a room-temperature at of 18 s and of 2 s [27]. These properties are particularly remarkable since both boron and nitrogen exhibit isotopically-abundant nuclear spins ( = 1, = 3, and = 3/2). Suppression of the nuclear spin bath using a second microwave source centred on a hyperfine line was found to improve and up to 25 s and 7.5 s, respectively. Work continues optimising the synthesis conditions for hBN to enhance the quantum properties of the defects. For instance, the contrast of -centres can be increased to 10 at room temperature by controlling the material purity and electron irradiation dose, though appears to remain phonon limited [111].
@@
Due to hBN’s potential as a tunable single photon source, significant effort has been made to control carbon-defect structures through careful synthesis. CVD-grown hBN has been found to contain a plethora of optically-addressable carbon-based spin species with stable ZFLs in the visible range; though most have yet to be conclusively structurally characterised, or even the multiplicity conclusively determined (see Figure 6d for candidate structures and Figure 6e for the spin initialisation scheme). For example, three distinct so called ""-defects, likely stemming from the substitution of a boron or nitrogen atom by carbon ("D1", "D2", "D3", all likely S=1/2), demonstrate similar spin relaxation and coherence quantum properties (0.2 s) with a continuous wave (CW)-optical contrast up to 20 at 5 K [30] (Figure 6f). Their inhomogeneous linewidths are also limited by the nuclear spin bath (40-60 ns), where local spins also bestow an apparent zero-field splitting of between 8-10 MHz.
@@
Though ensemble measurements remain challenging due to carbon-doped hBN’s broad emission profile, a recent report has shown that it is possible to survey over 400 defect centres in CVD-grown samples. Collectively known as ""-centres, careful measurements of individual spin-centres reveal distinctive triplet character with a ZFS of 15 MHz in contrast to most other accepted carbon-defects found in hBN [112]. The triplet identity was later unambiguously characterised and confirmed for a series of defects with significantly larger ZFS with |D|2 GHz and |E|50-80 MHz [29] (Figure 6g and h). The room-temperature was measured to be 20030 s, 1 s and 100 ns, competitive with -triplet centres. However, these properties are matched with a remarkable ODMR contrast of almost 50 (though it decreases with an applied magnetic field), giving rise to projected D.C. magnetic field sensitivity of up to 3 T Hz-1/2. A recent survey of similar defects demonstrated that under high microwave and optical pump powers, the optical contrast could reach as high as 90, though most exhibit <50 despite similar ZFS, and [113]. The authors explore the potential of this material as a medium for D.C. vectorial quantum magnetometry by taking advantage of the defects’ low symmetry and in-plane quantisation axis. Though yet-to-be experimentally verified, the authors hypothesise that applying a field bias along a different material axis corresponding to each of the three triplet states, Tx, Ty, and Tz, should result in subtle changes in contrast that can be related to particular Tn states to yield 3-vector components of the applied magnetic field.
@@
More complex carbon species demonstrating similar coherent behaviour at room temperature have been explored using metal-organic vapour-phase epitaxy with carbon-based additives [114, 28]. The so-called ""-defect (potential identity ) is another potential S=1/2 species or, two weakly coupled S=1/2 centres, has recently been reported as being co-located with -centres following electron irradiation of hBN flakes [28]. While their and are similar to -species (13 s and 80 ns, respectively), it can be used to "sense" -centres through spin cross relaxation at an appropriate applied magnetic field. Due to its out-of-plane magnetic quantisation axis (versus the in-plane axis of -centres), it can also more effectively be used to image magnetic structure on underlying substrates with an estimated DC magnetic sensitivity of 1 T Hz-1/2.
2.3 Molecular FOAMs
2.3.1 p-block molecular spin centres

### v3

@@ Abstract
I Introduction
@@
To be useful, spintronic materials must enable both single-spin addressability and long storage times at high levels of fidelity. These properties are, in turn, limited by the materials’ spin-lattice relaxation () and spin decoherence times (, also known as phase memory time). and represent the time it takes for the ”longitudinal” and ”transverse” magnitude of the spin state vector, respectively, to decay by a factor of e. In practical terms, is the lifetime of a non-Boltzmann population (Figure 1b), and, put simply by DiVincenzo [8], characterises the interactions of a qubit with its environment. These interactions are ultimately responsible for causing a qubit in a generic (pure) state to relax to a mixed state, leading to the loss of quantum information via phase decoherence (Figure 1c). These parameters can be measured using electron spin resonance (ESR) techniques, which measure light/microwave absorptions and emissions from a material as a function of frequency or magnetic field strength (Figure 2). is typically measured by Hahn-Echo pulses but can also be determined from the homogeneous emission linewidth of its resonance frequency, (Figure 1d). In practice, spin systems exhibit an inhomogeneously broadened linewidth, , due to differences in the magnetic field, B, at the locations of otherwise equivalent spins (see Figure 1e). These differences are in turn due to differences in the locations and orientations of the electronic and/or nuclear spins that surround each measured spin. Following Chavhan et al., the relationship between and can be represented as 1/ = 1/ + B, where is the gyromagnetic ratio [9]. In general, sets an upper limit on and, since energy relaxation proceeds more slowly than decoherence, is (often substantially) longer than [10]. Likewise, sets an upper limit on (the pulse sequence used to measure deliberately attempts to reduce the effects of inhomogeneity). Thus, .
@@
Quantum systems capable of operating at room temperature open up many additional applications that the overhead of cryogenic operation precludes. But, achieving high fidelity (in the initialisation, gate operations and read-out) of room-temperature qubits remains extremely challenging. In this review, we have attempted to survey the available ”fully optically addressable materials” (FOAMs) that are most directly relevant to quantum information and sensing technology [11]. We do not provide an exhaustive account of all reported spin systems. Rather, we present the different approaches, materials-wise, that have shown promise and report the properties (where available) of the best-performing representatives of each approach.
II Overview of Candidate Spin Qubit Materials
@@
FOAMs are an attractive option for quantum applications since they typically involve spin levels that are energetically separated by far more than and can be initialised using visible/near-infrared light to generate quasi-”pure” quantum states. Furthermore, optical signals, being themselves far larger than , are less affected by thermal noise and benefit from excellent single-photon detectors that are available even at room temperature. Since the energies of the photons involved in both processes greatly exceed , room-temperature operation at high levels of fidelity is possible. Such materials are usually probed using ODMR spectroscopy, where the quantum sensing sensitivity for a.c. magnetic fields (for example), , is proportional to , where is the duration of the readout process, is the measurement contrast, is the number of spins and is the number of scan averages. Relatively few FOAMs have been demonstrated so far and are often limited by either low photoluminescence yields or relatively short-lived spin polarizations () and coherences (), which have been summarised in Figure 3 to support our discussion. As a prerequisite for inclusion in Figure 3, we have opted to follow the criteria put forward by Weber et al., for the realization of an optically addressable solid-state spin qubit (that is useable)[12]:
@@
II.1 3D spin-defect FOAMs
II.1.1 Diamond spin centres
A well-established example of a FOAM which satisfies the above criteria is the negatively charged nitrogen-vacancy centres in diamond (herein simply, NV-diamond), which utilise the triplet ground state and hyperfine splitting arising from the NV-centre. Initialisation (i.e., electron spin polarisation) is achieved by optical excitation with green light. The newly generated excited states either relax by emission of 637 nm light or undergo intersystem crossing (ISC) into a metastable singlet state (Figure 4a). Repopulation of the ground state then follows an intermediate relaxation between two singlet states (with emission at 1042 nm) and finally spin-selective ISC into the T0 sublevel, resulting in a strong spin polarisation. At zero-applied magnetic field (ZF) the T±1 states are degenerate due to the defect’s C3V symmetry. The application of a magnetic field lifts this degeneracy, which enables spin manipulation using microwave pulses, thereby modifying the fluorescence at characteristic Zeeman splitting frequencies (), which can be detected following subsequent optical excitations and fluorescence back to T0.
NV-diamond has been a cornerstone of ODMR-based quantum sensing due to its robust spin properties even at room temperature. Due to the relatively low spin densities (resulting in small dipolar coupling between spins) and the mismatch between diamond lattice vibrations (phonons) and the Larmor frequency of electron spins in NV-’s S=1 ground state, NV-diamond can exhibit s of several milliseconds at room temperature [60, 61]. The nature and mechanism of its spin-lattice relaxation as a function of temperature, which includes an ‘Orbach-like” process (dependent on the phonon density at the spin transition frequency) and spin-phonon Raman scattering (of either first [62] or else second [61] order) scaling like , have been repeatedly investigated.
@@ II.1.1 Diamond spin centres
II.1.2 Silicon Carbide Spin Centres
@@
Beyond diamond, silicon carbide (SiC) also shows promising optical and coherence properties for quantum applications [85, 86, 87, 88, 89, 90]. SiC is a complex material with over 200 polymorphs. It is also used commercially in electronics and hence benefits from decades of manufacturing experience. Mercifully, the study of quantum systems has largely been restricted to 3C-, 4H-, and 6H-SiC, where C and H signify the cubic and hexagonal structures, respectively, and the preceding number designates its polytype [86] (see Figure 5a). Pure SiC has a wide bandgap (2 - 3 eV), weak spin-orbit coupling, and a naturally low abundance of nuclear spins. Importantly, it is capable of harbouring several different optically addressable colour centres with (often) near-infrared emission and record ODMR contrasts[31]. The most commonly studied defects consist of neutral or negatively-charged monovacancies (VC or VSi) and divacancies (VV) but can also include carbon anti-site vacancies (CSiVC) (see Figure 5b and c), charged NVs, Cr4+- [41, 91], V3+, V4+ [92], Mo5+ [44], and Ti-centres [93, 86, 90]. Further layers of complexity are added by consideration of additional oxidation states and crystallographic sites within each polytype. Understandably, studies often focus on the ”best performing” defect within a given sample, and seldom are single samples of SiC homogeneous. This represents the material’s most significant practical limitation as a spin qubit candidate compared to other platforms.
@@
Nevertheless, several defects exhibit outstanding spin-optical properties, and the fabrication challenges are beginning to be addressed. Negatively charged VSi in isotopically purified (28Si) 4H-SiC exhibits the largest ODMR contrast at 97 at 4 K [31]. With an S=3/2 ground state, its associated ZFS is a few MHz with degenerate pairs of —1/2 and —3/2 states (Figure 5d). Under a magnetic field (82 mT) precisely aligned to the crystallographic c-axis, this degeneracy is lost and in the excited, so-called ”V1” state, —3/2 shifts higher energy than the —1/2 states. To achieve 97 contrast, the authors first equilibrate the spin populations using a 40 s off-resonance pump (at 730 nm), followed by an on-resonance pump (at 861 nm) lasting up to 80 s. On-resonance optical pumping causes —3/2 states to selectively decay into a non-radiative metastable state, followed by spin-selective repopulation of —1/2 ground states. This leads to a spin polarisation of up to 90 which is paired with record s and s of 0.80.12 ms and 302 s, respectively. More recently, it was shown that VSi defects can be implanted into nanophotonic waveguides fabricated from 4H-SiC while also controlling the alignment of individual defects and maintaining excellent spin-optical properties [32]. Here, of bulk VSi-centres was measured at 344 s at 10 K, whilst those in the 1 m diameter waveguides measured at 9.40.7 s. The coherence properties can be further improved by a factor of 10 through a regime of isotopic purification, and another factor of 5 by reducing strain inhomogeneity through a regime of annealing. Using this combined approach, Lekavicius et al., enhanced from 400 ns to 20 s at room-temperature [33].
@@
II.1.3 Inorganic Semiconductor Spin Centres
Beyond vacancy systems, several metal-ion centres have demonstrated interesting spin-optical properties. By far the best, however, is Cr4+ in 4H-SiC. As a S=1 species, this material exhibits ¿1 second, and and = 81 s and 317 ns, respectively, at 15 K [91]. Importantly, it also demonstrates a 79 contrast, marking it as a system with some of the highest optical readout fidelity. The spin-optical properties are markedly impaired in GaN host, which demonstrates 27x broader emission linewidths due to interactions of Cr4+-centres with the surrounding spin bath [41].
II.2 2D- and van der Waals FOAMs
@@ II.1.3 Inorganic Semiconductor Spin Centres
The prototypical vdW material, graphene, was an early candidate for harbouring spin qubits; however, the metallic band structure of graphene is not conducive to forming luminescent spin centres. Instead, its isoelectronic analogue, hBN, has been the centre of remarkable advancement as a 2D-optically addressable platform at room temperature, building off its success as a single-photon source [103]. Magnetically-active spin centres in hBN were identified by EPR in the 1960s, with hyperfine structures giving clues to their identity as S=1/2 boron centres and carbon impurities [104, 105, 106]. However, the first optically addressable spin centres were only identified recently as triplets stemming from a negatively charged vacancy, likely originating from a missing boron atom, so-called -centres [107, 108, 109, 110] (Figure 6a). Similarly to NV-diamond centres, optical initialisation is achieved by photoexcitation with 532 nm light, leading to ISC into a metastable singlet, which spin-selectivity repopulates the fluorescent ms=0 state (TZ at zero-applied field, Figure 6b). The readout of the spin state occurs through 850 nm fluorescence with pulsed ODMR contrast of 1 at room temperature and almost 30 at 10 K, revealing a ZFS of —D—3.5GHz (Figure 6c). Spin-relaxation and Hahn-echo measurements reveal a room-temperature at of 18 s and of 2 s [27]. These properties are particularly remarkable since both boron and nitrogen exhibit isotopically-abundant nuclear spins ( = 1, = 3, and = 3/2). Suppression of the nuclear spin bath using a second microwave source centred on a hyperfine line was found to improve and up to 25 s and 7.5 s, respectively. Work continues optimising the synthesis conditions for hBN to enhance the quantum properties of the defects. For instance, the contrast of -centres can be increased to 10 at room temperature by controlling the material purity and electron irradiation dose, though appears to remain phonon limited [111].
@@
Due to hBN’s potential as a tunable single photon source, significant effort has been made to control carbon-defect structures through careful synthesis. CVD-grown hBN has been found to contain a plethora of optically-addressable carbon-based spin species with stable ZFLs in the visible range; though most have yet to be conclusively structurally characterised, or even the multiplicity conclusively determined (see Figure 6d for candidate structures and Figure 6e for the spin initialisation scheme). For example, three distinct so called ””-defects, likely stemming from the substitution of a boron or nitrogen atom by carbon (”D1”, ”D2”, ”D3”, all likely S=1/2), demonstrate similar spin relaxation and coherence quantum properties (0.2 s) with a continuous wave (CW)-optical contrast up to 20 at 5 K [30] (Figure 6f). Their inhomogeneous linewidths are also limited by the nuclear spin bath (40-60 ns), where local spins also bestow an apparent zero-field splitting of between 8-10 MHz.
@@
Though ensemble measurements remain challenging due to carbon-doped hBN’s broad emission profile, a recent report has shown that it is possible to survey over 400 defect centres in CVD-grown samples. Collectively known as ””-centres, careful measurements of individual spin-centres reveal distinctive triplet character with a ZFS of 15 MHz in contrast to most other accepted carbon-defects found in hBN [112]. The triplet identity was later unambiguously characterised and confirmed for a series of defects with significantly larger ZFS with —D—2 GHz and —E—50-80 MHz [29] (Figure 6g and h). The room-temperature was measured to be 20030 s, 1 s and 100 ns, competitive with -triplet centres. However, these properties are matched with a remarkable ODMR contrast of almost 50 (though it decreases with an applied magnetic field), giving rise to projected D.C. magnetic field sensitivity of up to 3 T Hz-1/2. A recent survey of similar defects demonstrated that under high microwave and optical pump powers, the optical contrast could reach as high as 90, though most exhibit ¡50 despite similar ZFS, and [113]. The authors explore the potential of this material as a medium for D.C. vectorial quantum magnetometry by taking advantage of the defects’ low symmetry and in-plane quantisation axis. Though yet-to-be experimentally verified, the authors hypothesise that applying a field bias along a different material axis corresponding to each of the three triplet states, Tx, Ty, and Tz, should result in subtle changes in contrast that can be related to particular Tn states to yield 3-vector components of the applied magnetic field.
@@
More complex carbon species demonstrating similar coherent behaviour at room temperature have been explored using metal-organic vapour-phase epitaxy with carbon-based additives [114, 28]. The so-called ””-defect (potential identity ) is another potential S=1/2 species or, two weakly coupled S=1/2 centres, has recently been reported as being co-located with -centres following electron irradiation of hBN flakes [28]. While their and are similar to -species (13 s and 80 ns, respectively), it can be used to ”sense” -centres through spin cross relaxation at an appropriate applied magnetic field. Due to its out-of-plane magnetic quantisation axis (versus the in-plane axis of -centres), it can also more effectively be used to image magnetic structure on underlying substrates with an estimated DC magnetic sensitivity of 1 T Hz-1/2.
II.3 Molecular FOAMs
II.3.1 p-block molecular spin centres